OpenStudy (anonymous):

ATT Myininaya: The gravitational force, F, on a rocket at a distance, r, from the center of the earth is given by F=k/r^2 where k = 10^13 newton • km2. When the rocket is 10^4 km from the center of the earth, it is moving away at 0.2 km/sec. How fast is the gravitational force changing at that moment? Give units. (A newton is a unit of force.)

myininaya (myininaya):

so it wants to know how fast F is changing so we need to find F'

myininaya (myininaya):

is this calculus or physics?

OpenStudy (anonymous):

Calculus

myininaya (myininaya):

it looks like k is a constant

OpenStudy (anonymous):

oh maybe

OpenStudy (anonymous):

that is where I may have gone wrong

OpenStudy (anonymous):

This is a related rates problem

OpenStudy (anonymous):

you may have to do df/dt

myininaya (myininaya):

\[F'=\frac{0 \cdot r^2-k \cdot 2rr'}{r^4}=\frac{-2kr'}{r^3}\]

myininaya (myininaya):

so we are given everything

OpenStudy (anonymous):

huh?

myininaya (myininaya):

i found F'

OpenStudy (anonymous):

I got a diff answer

OpenStudy (anonymous):

F ' =-2k/r^3

OpenStudy (anonymous):

i used the product rule

myininaya (myininaya):

\[(r^2)'=2r^{2-1}r'=2r^1r'=2rr'\]

myininaya (myininaya):

we can use the product rule if you want

OpenStudy (anonymous):

I just used the quotient rule and still got a diff answer

myininaya (myininaya):

\[F=kr^{-2} > F'=(-2)kr^{-2-1}r'=-2kr^{-3}r'=\frac{-2kr'}{r^3}\]

OpenStudy (anonymous):

what is r'?

myininaya (myininaya):

r is a function of time

OpenStudy (anonymous):

I am a drop confused

myininaya (myininaya):

r' is the rate that the distance is changing from the center of the earth

myininaya (myininaya):

\[[r(t)]'=r'(t)\]

OpenStudy (anonymous):

oh it is dr/dt?

OpenStudy (anonymous):

We didn't derive it normally we are doing like the related rate format

myininaya (myininaya):

these mean the same thing \[r'(t)=\frac{dr}{dt}\]

myininaya (myininaya):

just different notation

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

ya i was just getting confused a lil

myininaya (myininaya):

one is i think newton's notation and the other is lebniz notation i believe

myininaya (myininaya):

but they mean the samething

myininaya (myininaya):

\[F'(t)=[k \cdot [r(t)]^{-2}]'=k[[r(t)]^{-2}]'=k(-2)[r(t)]^{-2-1}r'(t)=-2kr'(t)[r(t)]^{-3}\]

OpenStudy (anonymous):

oh. i get it. I guess my book only shows liebniz's notation

myininaya (myininaya):

maybe i usually use newton's notation but that doesn't mean lebniz's is inferior notation

OpenStudy (anonymous):

ya i like newton's notation better but my book shows me this way df/dx ...

OpenStudy (anonymous):

is df/dt=-2(10^9)*.2?

myininaya (myininaya):

\[r=10^4 km ; r'=.2 km/\sec ; k=10^{13} newton \cdot km2\]

OpenStudy (anonymous):

oh whoops i made a mistake

OpenStudy (anonymous):

i forgot the cubed in the denomanator

OpenStudy (anonymous):

df/dt=2(10)*.2

OpenStudy (anonymous):

which is 8?

myininaya (myininaya):

\[F'=\frac{-2 k r'}{r^3}=\frac{-2(10^{13} N/km^2) \cdot (0.2 km/\sec)}{(10^4 km)^3}\] \[=\frac{-2 (10^{13}N/km^2) \cdot (\frac{2}{10} km /\sec)}{10^{4 \cdot 3} km^3}\] \[=\frac{-4 (10^{12} N/(km \cdot \sec))}{10^{12 }km^3}\] \[=-4 \frac{N}{km \cdot \sec} \cdot \frac{1}{km^3}=-4 \frac{N}{km^4 \sec}\] what did i miss?

OpenStudy (anonymous):

i think I went wrong somewhere but I will have to recheck it

myininaya (myininaya):

and maybe r' should have been negative since is says moving away

myininaya (myininaya):

and therefore F'= +4 not -4

OpenStudy (anonymous):

I may have punched accidentally the wrong number into the calculator and that is y i got the wrong number

OpenStudy (anonymous):

I think it is positive

myininaya (myininaya):

yes i think so too

OpenStudy (anonymous):

since the further away it is the larger r is

myininaya (myininaya):

oh wait are you talking about F' because thats what i mean

myininaya (myininaya):

r' is negative F' is positive

OpenStudy (anonymous):

yes that is what I meant

OpenStudy (anonymous):

meaning it shldn't be -.2

OpenStudy (anonymous):

Thanks for ur help

OpenStudy (anonymous):

Most ppl on here can't answer these questions. There are only a few ppl and hey r not on at all times

myininaya (myininaya):

now that i have read it couple of times r' is increasing not decreasing so r' is positive so F' is negative

OpenStudy (anonymous):

y is that?

OpenStudy (anonymous):

meaning that r'=-.2?

myininaya (myininaya):

no increasing would mean r'=.2

myininaya (myininaya):

decreasing means r'=-.2

OpenStudy (anonymous):

right so that is what we have right now

myininaya (myininaya):

but the distance is getting greater so r'=positive .2

OpenStudy (anonymous):

ok

myininaya (myininaya):

yes thats how i solved the problem initially

OpenStudy (anonymous):

Thanks;-D

myininaya (myininaya):

F'=-4 this means the gravitationial force is decreasing at a rate of 4 N/km^4 per sec

myininaya (myininaya):

if you wanted to know how to interpret that

OpenStudy (anonymous):

I know u r gonna kill me but isn't it km^3

OpenStudy (anonymous):

oh i see it

myininaya (myininaya):

\[F'=\frac{-2 k r'}{r^3}=\frac{-2(10^{13} N/km^2) \cdot (0.2 km/\sec)}{(10^4 km)^3} \] if i were to mess with the units i would have \[\frac{\frac{N}{km^2} \cdot \frac{km}{\sec}}{km^3}=\frac{\frac{N}{km \cdot \sec}}{km^3}\] \[=\frac{N}{km \cdot \sec} \cdot \frac{1}{km^3}=\frac{N}{km^4 \cdot \sec}\]

OpenStudy (anonymous):

OMG i wld not be able to solve that on my own!!!! I wld have totally left that out. I don't know how to solve wordproblems properly

myininaya (myininaya):

i usually don't bother with the units

myininaya (myininaya):

but they are just like numbers all you have to is treat them like numbers i mean

OpenStudy (anonymous):

well we had to write it in u

OpenStudy (anonymous):

units. Ok thanks for ur help. I really appreciate it. What level math r u in?

OpenStudy (anonymous):

oh wow good luck

myininaya (myininaya):

i need to go back and read some stuff

myininaya (myininaya):

i have alot of work to do lol

OpenStudy (anonymous):

y didn't u go into math u really good?

myininaya (myininaya):

i got my masters in mathematics

myininaya (myininaya):

i think thats enough

OpenStudy (anonymous):

ya I am aiming for that to. I am just in the beggining

myininaya (myininaya):

you will get there and you will probably be tons smarter than me by then

OpenStudy (anonymous):

If i ever need help will u be able to help me?

myininaya (myininaya):

if i'm around, i would be happy to help

OpenStudy (anonymous):

Oh my u sound like my dad already!!!

OpenStudy (anonymous):

Nice to meet you

myininaya (myininaya):

i hope i will never be anyone's dad since i'm a chick lol